OpenStudy (goldeverything):

Solve negative two v plus thirteen equals negative seven for v. negative ten negative three three ten MEDAL WILL BE GIVEN

OpenStudy (campbell_st):

so it looks like -2v + 13 = -7 subtract 13 from both sides of the equation then divide the by 2 for the answer.

OpenStudy (anonymous):

\[-2v + 13 = -7 \] -2v + 13 = -7 -13 -13 ----- ----- -2v = -20 _____ ____ -2 -2 v = ____